### v1

# Efficient Multivariate Initial Sequence Estimators for MCMC

Abstract
Estimating Monte Carlo error is critical to valid simulation results in Markov chain Monte Carlo (MCMC) and initial sequence estimators were one of the first methods introduced for this. Over the last few years, focus has been on multivariate assessment of simulation error, and many multivariate generalizations of univariate methods have been developed. The multivariate initial sequence estimator is known to exhibit superior finite-sample performance compared to its competitors. However, the multivariate initial sequence estimator can be prohibitively slow, limiting its widespread use. We provide an efficient alternative to the multivariate initial sequence estimator that inherits both its asymptotic properties as well as the finite-sample superior performance. The effectiveness of the proposed estimator is shown via some MCMC example implementations. Further, we also present univariate and multivariate initial sequence estimators for when parallel MCMC chains are run and demonstrate their effectiveness over popular alternative.
Keywords— Autocovariances, covariance matrix estimation, central limit theorem, parallel Markov chains
1 Introduction
Variance assessment for Markov chain Monte Carlo (MCMC) estimators is critical to accurate Bayesian inference. In MCMC, given a target distribution $\pi$ defined on the probability space $(\mathcal{X},\mathcal{B}(\mathcal{X}))$, an ergodic Markov chain $\{X_{t}\}_{t\geq 1}$ is constructed such that $\pi$ is its stationary distribution. When interest is in $\theta=\mathbf{E}_{\pi}(g(X))$ for a function $g:\mathcal{X}\to\mathbb{R}^{d}$, Birkhoff’s ergodic theorem allows for the following estimator to be strongly consistent:
$\bar{g}_{n}:=\dfrac{1}{n}\sum_{t=1}^{n}g(X_{t})\,.$
Over the years, many output analysis tools have focused on assessing Monte Carlo error in $\bar{g}_{n}$. This can be achieved if $g$ and the Markov chain exhibit a central limit theorem (CLT) such that there exists a $d\times d$ positive-definite covariance matrix $\Sigma$ so that as $n\to\infty$,
[…]
Estimating $\Sigma$ is critical for output analysis as it represents the variability in the estimation of $\theta$. Let $\zeta_{i}$ denote the lag-$i$ auto-covariance matrix
$\zeta_{i}=\mathbf{E}_{\pi}\left[\left(g(X_{1})-\theta\right)\left(g(X_{i+1})-\theta\right)^{\top}\right]\,.$
Then, when a CLT holds for $\bar{g}_{n}$, $\Sigma=\sum_{i=-\infty}^{\infty}\zeta_{i}$. It is naturally challenging to estimate $\Sigma$ and many efforts have been made in this direction, both for univariate $(d=1)$ and multivariate problems. Seila, (1982) proposed regeneration-based estimators that exhibit good properties but are challenging to employ in practice and require studying individual Markov chain kernels. As a consequence, estimators that can be implemented without knowledge of the kernel are more favored. These include batch-means estimators, spectral variance estimators, and initial sequence estimators (ISE). Amongst these, batch-means (Chen and Seila, , 1987; Jones et al., , 2006; Vats et al., , 2019) are computationally efficient but can exhibit significant undesirable underestimation (Vats and Flegal, , 2021). Spectral variance estimators (Andrews, , 1991; Flegal and Jones, , 2010; Vats et al., , 2018), can control this underestimation to some extent but are more computationally involved.
As pointed out in Vats and Flegal, (2021), in the presence of positive auto-correlation in the Markov chain, underestimation in $\Sigma$ can greatly impact the quality of inference. Specifically, underestimating the asymptotic variance leads to inflated effective sample size estimates, early termination of the MCMC, and a false sense of security about the quality of estimation. For reversible Markov chains, univariate initial sequence estimators (ISE), first proposed by Geyer, (1992), are known to be asymptotically conservative, exhibiting reduced underestimation in finite samples. ISEs are formed by truncating the sum of the sample auto-covariances, where the truncation point is data-dependent, and is sequentially determined. For the one-dimensional case, all sample auto-covariances can be obtained quickly using a fast Fourier transform (FFT) with a cost of $\mathcal{O}(n\log n)$, alleviating the burden of sequentially calculating sample autocovariances. Geyer, ’s (1992) ISE has over the years been studied and employed in many applications. Fabreti and Höhna, (2022) demonstrated its superior performance for analyzing the quality of phylogenetic MCMC, and Ardia and Bluteau, (2017); Carpenter et al., (2017); Geyer et al., (2015) employ the ISE estimator as default in their MCMC software.
In most MCMC applications $g$ is multivariate and marginal assessment of Monte Carlo error is insufficient to capture the complex inter-dependence across the components of $\bar{g}_{n}$. Similar to batch-means and spectral variance estimators, multivariate extensions of Geyer, ’s (1992) ISE have been proposed by Dai and Jones, (2017); Kosorok, (2000). However, multivariate ISEs no longer benefit from fast implementation via FFTs and require repeated determinant calculations to obtain the truncation point. The resulting estimator is prohibitively slow.
We propose a new multivariate version of the ISE estimator which combines the finite-sample performance of the ISE with the computational efficiency of batch-means estimators while retaining the asymptotic theoretical properties of the multivariate estimator of Dai and Jones, (2017). Section 3 elaborates on the proposed estimator, its computational cost, and its theoretical properties. We acknowledge that running parallel Markov chains is commonplace in MCMC. However, univariate or multivariate ISEs have not been developed for this framework. In Section 5 we present a parallel-chain version of Geyer, ’s (1992) ISE and then adapt it for a computationally efficient multivariate estimator. We show that the parallel-chain version of the ISE estimator retains the asymptotic properties of the single-chain estimator. Our examples in Section 4 highlight the improved estimation quality and the drastically reduced computation time for our method compared to that of Dai and Jones, (2017). We first compare the performance of our proposed estimator to the existing ones for a toy problem where the true value of $\Sigma$ is known. Our second example is that of an MCMC employed in a Bayesian Poisson regression model analyzing spike train data [^1] Code is available at https://github.com/Arkagit/Efficient-Initial-Sequence-estimator. In both our implementations it is evident that our proposed estimator, in addition to being significantly faster than current multivariate ISE, exhibits better finite-sample properties.
2 Initial Sequence Estimators
2.1 Univariate Initial Sequence Estimators
We present the original positive ISE of Geyer, (1992) for the univariate case. For $d=1$, if $\gamma_{i}$ denote the lag-$i$ autocovariance,
$\gamma_{i}=\text{Cov}_{\pi}(g(X_{1}),g(X_{i+1}))\,,$
then the asymptotic variance in the Markov chain CLT in (1) can be expressed as
$\sigma^{2}=-\gamma_{0}+2\sum_{i=0}^{\infty}\left(\gamma_{2i}+\gamma_{2i+1}\right)\,.$
For $k=0,\dots,n-1$, denote the sample autocovariance as
$\hat{\gamma}_{i;n}:=\dfrac{1}{n}\sum_{t=1}^{n-i}(g(X_{t})-\bar{g}_{n})(g(X_{t+i})-\bar{g}_{n})\,.$
Ideally, we would like to estimate $\sigma^{2}$ using the sum of all $\hat{\gamma}_{i;n}$’s, however, this leads to an inconsistent estimator. Define $\Gamma_{i}:=(\gamma_{2i}+\gamma_{2i+1})$ for all $i\geq 0$ and let $\hat{\Gamma}_{i;n}=(\hat{\gamma}_{2i;n}+\hat{\gamma}_{2i+1;n})$. For reversible Markov chains, Geyer, (1992) showed that $\Gamma_{i}>0$ for all $i$. This result is then used to truncate the sum of the $\hat{\gamma}_{i;n}$, yielding the ISE. That is,
$\hat{\sigma}^{2}_{\text{ISE}}=-\hat{\gamma}_{0;n}+2\sum_{i=0}^{k_{n}}\hat{\Gamma}_{i;n}\,,$
where $k_{n}$ is the largest integer such that $\hat{\Gamma}_{i+1;n}>\hat{\Gamma}_{i;n}$ for all $i\in\{1,2,\ldots,k_{n}\}$. That is, the sum of sample auto-covariances is stopped as soon as there is evidence that the auto-covariance for that lag is not well estimated.
Theorem 1.
(Geyer, , 1992, Theorem 3.2) For a $\pi$-reversible, Harris ergodic Markov chain such that a univariate CLT holds for $g$,
$\liminf_{n\rightarrow\infty}\hat{\sigma}^{2}_{\text{ISE}}\geq\sigma^{2}\qquad{\text{with probability $1$}}\,.$
Theorem 1 establishes asymptotic conservativeness of the $\hat{\sigma}^{2}_{\text{ISE}}$, and numerical experiments through the years have also established good finite-sample properties. Particularly for slowly mixing Markov chains for which truncation time $k_{n}$ can be large, implementation of the ISE is greatly benefited by an FFT which allows for a calculation of all $\hat{\gamma}_{i:n},i=0,1,2,\dots,n-1$ in $\mathcal{O}(n\log n)$ time. On the flip side, the strong consistency of the ISE has not been shown yet and remains an open question. Berg and Song, (2023) recently proposed a shape-constrained estimator of the sequence of $\gamma_{i}$, which they then used to obtain a moment least squares estimator (MLS) of $\sigma^{2}$ and obtained conditions for strong consistency. The computational complexity of the estimator is not clear, but in our simulations, it is evident that the MLS estimator requires more time than $\hat{\sigma}^{2}_{\text{ISE}}$.
2.2 Multivariate Initial Sequence Estimators
MCMC samplers are usually employed when $\pi$ is multidimensional and exhibits complex dependence across components. Consequently, multivariate error assessment in $\bar{g}_{n}$ is critical to MCMC output analysis. Kosorok, (2000) first extended Geyer, ’s (1992) univariate ISE to a multivariate case, and an improved alternative was presented by Dai and Jones, (2017). Although built in the same spirit as the univariate ISE, the practical limitations of the multivariate initial sequence estimators (mISE) are significant. Let $\hat{\zeta}_{i;n}$ denote the lag-$i$ sample auto-covariance matrix for $i=0,1,\ldots,(n-1)$
$\hat{\zeta}_{i;n}=\hat{\zeta}_{-i;n}^{\top}=\frac{1}{n}\sum_{t=1}^{n-i}(g(X_{t})-\bar{g}_{n})(g(X_{t+i})-\bar{g}_{n})^{\top}.$
Let $\hat{Z}_{i;n}=\left(\hat{\zeta}_{2i;n}+\hat{\zeta}_{2i+1;n}\right)$ and for $0\leq m\leq\lfloor n/2-1\rfloor$, define
$\hat{\Sigma}_{m;n}=-\hat{\zeta}_{0;n}+2\sum_{i=0}^{m}\hat{Z}_{i;n}.$
Here, $\hat{\Sigma}_{m;n}$ is the sum of sample auto-covariances up to the lag $2m$. Dai and Jones, (2017) showed that the population quantity $Z_{i}=\xi_{2i}+\xi_{2i+1}$ is positive-definite for all $i$ and further $Z_{i}-Z_{i-1}$ is also positive-definite for all $i\geq 1$. Using these properties, Dai and Jones, (2017) define their mISE as
$\hat{\Sigma}_{\text{ISE}}:=\hat{\Sigma}_{t_{n};n}$
where $t_{n}$ is the truncation time as described in Algorithm 1.
Data: $\{g(X_{1}),\ldots,g(X_{n})\}$
$\hat{\Sigma}\leftarrow-\hat{\zeta}_{0;n}$; $\triangleright$ Time Cost: $\mathcal{O}(nd^{2})$
for $i=0,1,\ldots,\lfloor(n-1)/2\rfloor$ do
       $\hat{\Sigma}=\hat{\Sigma}+2\hat{Z}_{i;n}$;
       if $\hat{\Sigma}$ is positive-definite then
             $s_{n}=i$; break;
       $\triangleright$ Time Cost: $\mathcal{O}(ns_{n}d^{2})$
for $i=(s_{n}+1),\ldots,\lfloor n/2-1\rfloor$ do
       $S=\hat{\Sigma}$;
       $\hat{\Sigma}=\hat{\Sigma}+2\hat{Z}_{i;n}$;
       if det($\hat{\Sigma}$) $\leq$ det($S$) then
             $t_{n}=(i-1)$; $\triangleright$ The truncation time
             $\hat{\Sigma}_{\text{ISE}}\leftarrow S$; break;
       $\triangleright$ Time Cost: $\mathcal{O}(nt_{n}d^{2})$
return $\hat{\Sigma}_{\text{ISE}}$
Algorithm 1 mISE of Dai and Jones, (2017)
Both the univariate ISE and mISE do not require any tuning. However, for slow mixing Markov chains, $t_{n}$ can be quite large and further, as $n$ increases, the quality of estimation of $Z_{i}$ improves, increasing the truncation time $t_{n}$. The following proposition establishes this phenomenon concretely.
Proposition 1.
Let $t_{n}$ be the truncation time for the mISE in Algorithm 1. Then with probability $1$
$t_{n}\rightarrow\infty\quad\text{ as }n\rightarrow\infty.$ (2)
Proof.
See Appendix A.1. ∎
Proposition 1 implies that for long-run MCMC samplers, the truncation time can be large. Algorithm 1 indicates the mISE estimator has an overall cost of $\mathcal{O}(nt_{n})$. Although the rate of increase of $t_{n}$ is not known, in Section 4, we empirically show that $t_{n}$ dominate $\log(n)$. Further, Algorithm 1 requires repeated and sequential calculation of determinants and thus is not parallelizable. Consequently, the mISE of Dai and Jones, (2017) can be prohibitively slow impeding its widespread use.
3 Covariance-Correlation ISE
The central idea for the proposed estimator is based on the observation that any covariance matrix can be decomposed to the product of a diagonal matrix of marginal standard deviations and its corresponding correlation matrix. That is, $\Sigma$ can be written as
$\Sigma=L\mathcal{R}L^{\top}$ (3)
where $\mathcal{R}$ is the correlation matrix corresponding to $\Sigma$ and $L$ is a diagonal matrix of marginal standard deviations. We will adopt different methods for estimating $L$ and $\mathcal{R}$. In order to do this, we first present the multivariate batch-means estimator.
For a Monte Carlo sample size of $n=a_{n}b_{n}$, let $b_{n}$ be the batch-size and $a_{n}$ be the number of batches. For $k=1,2,\dots,a_{n}$, denote the mean vector in each batch by $\tilde{g}_{k}:=b_{n}^{-1}\sum_{t=(k-1)b_{n}+1}^{kb_{n}}g(X_{t})$. The batch-means estimator of $\Sigma$ first presented by Chen and Seila, (1987) is defined as
$\hat{\Sigma}_{\text{BM}}:=\frac{b_{n}}{a_{n}-1}\sum_{k=1}^{a_{n}}(\tilde{g}_{k}-\bar{g}_{n})(\tilde{g}_{k}-\bar{g}_{n})^{\top}.$ (4)
Let $\left(\hat{\sigma}_{\text{BM}}^{(s)}\right)^{2}$ denote the $s$th diagonal element of $\hat{\Sigma}_{\text{BM}}$ and set
$\hat{L}_{\text{BM}}=\text{diagonal}\left(\hat{\sigma}_{\text{BM}}^{(1)},\hat{\sigma}_{\text{BM}}^{(2)},\dots,\hat{\sigma}_{\text{BM}}^{(d)}\right)\,.$ (5)
Using (4) and (5), we can obtain a batch-means estimator of the correlation matrix,
^$\hat{\mathcal{R}}_{\text{BM}}=\hat{L}_{\text{BM}}^{-1}\hat{\Sigma}_{\text{BM}}\hat{L}_{\text{BM}}^{-1}\,.$ (6)
In the presence of positive autocorrelation in the components of the process $\{g(X_{t})\}$, for any $s^{th}$ marginal, $\hat{\sigma}_{\text{BM}}^{(s)}$ underestimates $\sigma^{(s)}$ (Flegal and Jones, , 2010), with the underestimation being worse when autocorrelation is high. Underestimating these marginal variances has a significant impact on the quality of inference, however, bias in the estimation of the correlation matrix is more tolerable.
Let $\hat{L}_{\text{ISE}}$ denote the diagonal matrix of Geyer, ’s (1992) ISE standard deviations
$\hat{L}_{\text{ISE}}=\text{diagonal}\left(\hat{\sigma}_{\text{ISE}}^{(1)},\hat{\sigma}_{\text{ISE}}^{(2)},\dots,\hat{\sigma}_{\text{ISE}}^{(d)}\right)\,.$ (7)
Using ^$\hat{\mathcal{R}}_{\text{BM}}$ and $\hat{L}_{\text{ISE}}$ we obtain our proposed covariance-correlation ISE estimator of $\Sigma$
$\hat{\Sigma}_{\text{cc}}=\hat{L}_{\text{ISE}}\hat{\mathcal{R}}_{\text{BM}}\hat{L}_{\text{ISE}}\,.$ (8)
Since the marginal univariate ISEs are efficient, obtaining $\hat{L}_{\text{ISE}}$ is relatively cheap, and since batch-means estimators are fairly efficient, ^$\hat{\mathcal{R}}_{\text{BM}}$ also has low computational burden. This reduces the overall computational cost to $\mathcal{O}(n\log n)$, and in Section 4 we show that the computational cost is significantly lower than mISE. The marginal variances are well estimated by the ISE, and the resultant estimator is guaranteed to be positive semi-definite. Later in Section 4 we demonstrate these finite-sample properties of $\hat{\Sigma}_{\text{cc}}$ and here provide conditions under which $\hat{\Sigma}_{\text{cc}}$ is asymptotically conservative. This naturally requires conditions ensuring asymptotic properties of batch-means estimators.
Assumption 1.
[…]
$\left\|\sum_{t=1}^{n}g(X_{t})-n\theta-MB(n)\right\|<D\kappa(n)\,.$ (9)
There has been much work done in ensuring a strong invariance principle holds for MCMC (Jones et al., , 2006; Li and Qin, , 2024). Vats et al., (2018) show that $\kappa(n)=n^{1/2-\lambda}$ for some $20<\lambda<1/$ 2 for when $\{X_{t}\}$ is polynomially ergodic. Banerjee and Vats, (2024) show $\kappa(n)=n^{\beta}\log n$ for $\beta=\max\{1/(2+\delta),1/4\}$ for when the process is either geometrically or polynomially ergodic with moment conditions on $g$.
Assumption 2.
The batch-size $b_{n}$ should be chosen so that
  • 1.
    $b_{n}$ and $n/b_{n}$ are increasing integer sequences such that as $n\rightarrow\infty$, $b_{n}\rightarrow\infty$ and $n/b_{n}\rightarrow\infty$.
  • 2.
    there exists $c\geq 1$ such that $\sum_{n=1}^{\infty}(b_{n}/n)^{c}<\infty$.
  • 3.
    $b_{n}^{-1}\log(n)\kappa^{2}(n)\rightarrow 0$ as $n\rightarrow\infty$.
For the batch-means estimator to be strongly consistent, the number of batches and batch-size both must increase to infinity. Further, the batch-size is typically governed by the strength of the correlation in the Markov chain. It is common to use batch-sizes of the form $b_{n}=\lfloor n^{\nu}\rfloor$ for some $10<\nu<$ 1.
Theorem 2.
Let $\{X_{t}\}_{t\geq 1}$ be a $\pi$-reversible Harris ergodic Markov chain such that Assumption 1 holds for $g:\mathcal{X}\to\mathbb{R}^{d}$. If Assumption 2 holds, then we have
  1. (a)
    $\liminf_{n\rightarrow\infty}\text{det}\left(\hat{\Sigma}_{\text{cc}}\right)\geq\text{det}\left(\Sigma\right)\text{ with probability }1;\text{ and }$
  2. (b)
    for $\hat{a}_{ij}$ and $a_{ij}$ denoting the $(i,j)^{\text{th}}$ element of $\hat{\Sigma}_{cc}$ and $\Sigma$, respectively,
    $\liminf_{n\rightarrow\infty}|\hat{a}_{ij}|\geq|a_{ij}|\text{ with probability }1.$
Proof.
See Appendix A.2. ∎
Remark 1.
Proposition 2 presents convergence properties of the generalized variance of $\hat{\Sigma}_{\text{cc}}$ which is a similar result to that of mISE from Dai and Jones, (2017). The result implies that asymptotically, the confidence regions created from $\hat{\Sigma}_{\text{cc}}$ will not have a smaller volume than the oracle. Although generalized variance has its statistical importance, it does not comment on the estimation of all the cross-covariance terms of $\Sigma$. Part $(b)$ of the theorem establishes element-wise asymptotic conservativeness.
The univariate MLS estimator of Berg and Song, (2023) estimator can be modified similarly for the multivariate setting. We define $\hat{\Sigma}_{\text{MLS}}:=\hat{L}_{\text{MLS}}\hat{\mathcal{R}}_{\text{BM}}\hat{L}_{\text{MLS}}^{\top}$ where $\hat{L}_{\text{MLS}}=\text{diagonal}(\hat{\sigma}_{\text{MLS}}^{(i)};1\leq i\leq d)$. Under the conditions ensuring strong consistency of the MLS estimator from Berg and Song, (2023), $\hat{\Sigma}_{\text{MLS}}$ can also be shown to be strongly consistent. However, will show in Section 4 that this estimator although faster than the mISE, is significantly slower than $\hat{\Sigma}_{\text{cc}}$.
4 Numerical Implementations
4.1 Vector Autoregressive Process
We study the comparative performance of the estimators in a benchmark vector autoregressive (VAR) process example. The VAR example yields a setting where the true $\Sigma$ is known. Consider a VAR process of order $1$, $\{X_{t}\}_{t\geq 1}$, such that for a $d\times d$ matrix $\Phi$ and $t=1,2,\ldots$
$X_{t}=\Phi X_{t-1}+\epsilon_{t}\,,$
[…]
$\Sigma_{\text{True}}=(I_{d}-\Phi)^{-1}V+V(I_{d}-\Phi)^{-1}-V.$
We borrow the setup of Dai and Jones, (2017) and choose $\Omega=I_{d}$ and $\Phi=d^{-1}\mathcal{H}_{d}$ $\text{diag}(\rho^{-1},\rho^{-2},\ldots,\rho^{-d})\mathcal{H}_{d}^{\top}$ where $\mathcal{H}_{d}$ is the Hadamard matrix of dimension $d=12$. With this choice of $\Phi$ and $\Omega$, the process $\{X_{t}\}_{t\geq 1}$ is reversible. For $\rho=1.01$, we compare the performance of $\hat{\Sigma}_{\text{cc}}$ with the existing estimators. For $n=5\times 10^{3}\text{ to }5\times 10^{5}$, we assess the performance of estimators of $\Sigma$ by four metrics (i) estimated effective sample size (ESS), (ii) relative Frobenius norm, (iii) computation time, and (iv) coverage probability, all averaged over 1000 replications. Vats et al., (2019) define the estimated ESS as
$\widehat{\text{ESS}}=n\left(\frac{\text{det}(\hat{\zeta}_{0})}{\text{det}(\hat{\Sigma})}\right)^{1/d},$
where $\hat{\zeta}_{0}$ is the sample covariance matrix and $\hat{\Sigma}$ is the estimator of interest for $\Sigma$. Vats and Knudson, (2021) showed that there is a one-to-one correspondence between ESS and the $R$-hat of Gelman and Rubin, (1992). Consequently, stopping simulation when $\widehat{\text{ESS}}$ is larger than a pre-specified lower bound is standard, and thus it is critical to ensure estimators of $\Sigma$ do not underestimate $\text{det}(\Sigma)$. For $\|\cdot\|_{\text{F}}$ denoting Frobenius norm, the relative Frobenius norm is
Relative Frobenius $\text{Relative Frobenius Norm}=\frac{\|\Sigma-\hat{\Sigma}\|_{\text{F}}}{\|\Sigma\|_{\text{F}}}.$
Figure 1: (Top left) Relative Frobenius norm of estimators of $\Sigma$. (Top right) Estimated ESS$/n$. (Bottom left) Computation time in seconds. (Bottom right) Truncation time $t_{n}$ for mISE as compared to $\log n$.
We compare the $\hat{\Sigma}_{cc}$ (CC-ISE) with the Dai and Jones, ’s (2017) mISE estimator, along with batch-means (BM) and spectral variance estimator (SVE). Additionally, we also implement CC-MLS, the covariance-correlation estimator built with Berg and Song, ’s (2023) MLS estimator for the marginal variances. The results are provided in Figure 1. First, both the CC estimators yield the lowest relative Frobenius norm, significantly lower than the mISE estimator. For ESS, as explained earlier, in order to avoid early termination of simulation, underestimating the true quantity is more favorable than overestimating; here we see clearly that both the CC estimators converge to the true value from below. These gains in finite-sample performance are supported by quick computation. The mISE estimator is prohibitively slow, particularly for the large sample sizes typically found in MCMC. In contrast $\hat{\Sigma}_{\text{cc}}$ is much faster, only a little slower than SVE.
Over the 1000 replications, we also track the truncation time for the mISE estimator for varying sample sizes. In the bottom right plot of Figure 1, we demonstrate the rate of increase of $t_{n}$ for the mISE estimator as it compares to $\log(n)$. The plot shows that $t_{n}$ certainly dominates $\log(n)$ and diverges to infinity as chain length increases.
For assessing the coverage probability, we determine whether the asymptotic confidence ellipsoid contains the true value of $\mathbf{E}_{\pi}X=0$. An asymptotic $95\%$ confidence ellipsoid from (18) and using an estimator $\hat{\Sigma}$ of $\Sigma$ is
${C}_{0.95}=\{\mu\in\mathbb{R}^{d}:n(\bar{X}_{n}-\mu)^{\top}\hat{\Sigma}^{-1}(\bar{X}_{n}-\mu)<\chi^{2}_{d,0.95}\}.$
Over 1000 replications and for the competing estimators of $\Sigma$, we determine the number of times the estimated confidence ellipsoid contains 0. Results are in Table 1. As expected, BM and SVE have low coverage for smaller sample sizes due to the underestimation of $\Sigma$. Dai and Jones, ’s (2017) mISE also has low coverage at smaller sample sizes, indicating an underestimation of the target variance. On the other hand, CC-ISE exhibits improved coverage at low sample sizes as well as large sample sizes.
Method Chain Length
$5\times 10^{3}$ $10^{4}$ $5\times 10^{4}$ $10^{5}$ $5\times 10^{5}$
BM $0.474$ $0.664$ $0.883$ $0.887$ $0.952$
SVE $0.589$ $0.751$ $0.896$ $0.913$ $0.960$
ISE $0.651$ $0.778$ $0.906$ $0.913$ $0.960$
CC-ISE $0.715$ $0.883$ $0.948$ $0.962$ $0.974$
CC-MLS $0.623$ $0.824$ $0.936$ $0.944$ $0.971$
Table 1: Coverage probability at $95\%$ confidence at different chain lengths.
4.2 Bayesian Poisson Regression for Spike Train Data
Consider the spike train dataset from D’Angelo and Canale, (2023) which uses calcium imaging to observe neuron responses to external events. The data consists of $p=23$ factors as predictors for 920 responses of the count of activations in each neuron. D’Angelo and Canale, (2023) use the following Bayesian Poisson regression model to analyze this data.
For $i=1,2,\dots,920$, let $Y_{i}$ denote the random response and $y_{i}$ be the realized value of the count of the neuron activations. Further, let $\mathbf{x}_{i}\in\mathbb{R}^{23}$ denote the $i$th vector of predictors. For $\beta\in\mathbb{R}^{23}$, D’Angelo and Canale, (2023) propose the following model,
[…]
$\beta$ $\sim\text{Normal}(\mathbf{b},B).$
The posterior distribution for $\beta$ is unavailable and MCMC is employed to obtain posterior estimates. The Bayesian Poisson regression model is known to be challenging to sample from and D’Angelo and Canale, (2023) propose a novel Metropolis-Hastings proposal built specifically for this model. We employ their algorithm to generate samples from the posterior distribution. The Metropolis-Hastings sampler is available in the bpr R package (D’Angelo, , 2022).
Similar to D’Angelo and Canale, (2023), we set $\mathbf{b}=\mathbf{0}$ and $B=2I_{23}$. Over 100 replications, we run the MCMC algorithm for various Monte Carlo lengths and employ different estimators of $\Sigma$ to estimate the posterior mean vector. The true value of both $\bar{g}_{n}$ and $\Sigma$ is unknown here, so we rely on the estimation of ESS$/n$ and computation time to compare methods. The results are provided in Figure 2.
The average estimated ESS$/n$ is lowest for the CC-ISE estimator, although the correlation in the Markov chain is persistent enough to not allow for convergence from below. The mISE estimator has significantly larger ESS estimates. The mISE estimator comes with a high computational cost as well, taking as long as 5 minutes to compute the estimator. The CC-MLS estimator is also computationally involved, due to the repeated application of the MLS estimator for all 23 diagonals. Comparatively, the CC-ISE estimator is significantly cheaper to compute and yields desirable finite-sample performance.
Figure 2: (Left) Estimated ESS$/n$ for increasing chain lengths. (Right) Computational time for increasing chain lengths (in seconds).
5 Initial sequence estimators for parallel chains
Modern computers have multiple cores that allow for parallel runs of the MCMC to be easily run in the same clock time. Recently, Agarwal and Vats, (2022); Douc et al., (2022) and Gupta and Vats, (2020) demonstrated the utility of developing specialized methods for estimating $\Sigma$ in such a case. Agarwal and Vats, (2022) developed globally-centered auto-covariances to estimate $\zeta_{i}$, which we show can be easily adapted to construct a parallel chain version of Geyer, ’s (1992) ISE.
Let $\{X_{t}^{(m)}\}_{1\leq t\leq n}$ be the $m^{th}$ parallel $\pi$-reversible Markov chain for $m=1,\ldots,M$ generated using the same Markov transition kernel. Consider the case when $d=1$ and let the Monte Carlo mean for the $m$th chain be $\bar{g}^{(m)}:=n^{-1}\sum_{t=1}^{n}g(X_{t}^{(m)})$ and the global mean be $\bar{\bar{g}}:=M^{-1}\sum_{m=1}^{M}\bar{g}^{(m)}$. Agarwal and Vats, (2022) define the globally-centered lag-$i$ auto-covariance for the $m^{th}$ parallel chain to be
$\hat{\gamma}_{i;n}^{(m)}=\frac{1}{n}\sum_{t=1}^{n-i}\left(g(X_{t}^{(m)})-\bar{\bar{g}}\right)\left(g(X_{t+i}^{(m)})-\bar{\bar{g}}\right).$
The final estimator of the auto-covariance is
$\hat{\gamma}_{i;n}^{\text{G}}:=\dfrac{1}{M}\sum_{m=1}^{M}\hat{\gamma}_{i;n}^{(m)}\,.$ (10)
Agarwal and Vats, (2022) demonstrated how a small fix of global-centering can massively improve the estimation of auto-covariance, particularly for slow-mixing Markov chains or multi-modal targets. They specifically demonstrated how globally-centered auto-covariances when employed in spectral-variance estimators, lead to improved estimation of $\sigma^{2}$ (or $\Sigma$).
We propose a globally-centered adaptation of the initial sequence estimator of Geyer, (1992). Let $\hat{\Gamma}_{i;n}^{\text{G}}=(\hat{\gamma}^{\text{G}}_{2i;n}+\hat{\gamma}^{\text{G}}_{2i+1;n})$. We define the globally-centered ISE (G-ISE) as
$\hat{\sigma}^{2}_{\text{G-ISE}}=-\hat{\gamma}^{\text{G}}_{0;n}+2\sum_{i=0}^{k_{n}}\hat{\Gamma}^{\text{G}}_{i;n}\,,$ (11)
where $k_{n}$ is the largest integer such that $\hat{\Gamma}^{\text{G}}_{i+1;n}>\hat{\Gamma}^{\text{G}}_{i;n}$ for all $i\in\{1,2,\ldots,k_{n}\}$. The following theorem establishes the asymptotic conservation result of the GCC-ISE.
Theorem 3.
For $\pi$-reversible, Harris ergodic Markov chains such that a univariate CLT holds
$\liminf_{n\rightarrow\infty}\hat{\sigma}^{2}_{\text{G-ISE}}\geq\sigma^{2}\qquad\text{with probability $1$}\,.$
Proof.
See Appendix A.3. ∎
Having described the univariate globally-centered ISE estimator, we can now present the covariance-correlation globally-centered multivariate estimator. Let $\left(\hat{\sigma}_{\text{G-ISE}}^{(j)}\right)^{2}$ denote the globally-centered ISE estimator for the $j$th component of $\theta$. Let $\hat{L}_{\text{G-ISE}}$ denote the diagonal matrix of marginal G-ISE standard deviations
$\hat{L}_{\text{G-ISE}}=\text{diagonal}\left(\hat{\sigma}_{\text{G-ISE}}^{(1)},\hat{\sigma}_{\text{G-ISE}}^{(2)},\dots,\hat{\sigma}_{\text{G-ISE}}^{(d)}\right)\,.$ (12)
Gupta and Vats, (2020) present a globally-centered version of the batch-means estimator as well, which we denote by ^$\hat{\mathcal{R}}_{\text{G-BM}}$. Using ^$\hat{\mathcal{R}}_{\text{G-BM}}$ and $\hat{L}_{\text{G-ISE}}$, a globally-centered covariance-correlation ISE (GCC-ISE) of $\Sigma$ can be naturally constructed as:
$\hat{\Sigma}_{\text{G-cc}}=\hat{L}_{\text{G-ISE}}\hat{\mathcal{R}}_{\text{G-BM}}\hat{L}_{\text{G-ISE}}\,.$ (13)
Corollary 1.
Let $\{X_{t}^{(s)}\}_{t\geq 1}$ be $\pi$-reversible Harris ergodic Markov chains such that a CLT holds for a function $g$. If Assumptions 1 and 2 hold, we have
$\liminf_{n\rightarrow\infty}\text{det}\left(\hat{\Sigma}_{\text{G-cc}}\right)\geq\text{det}\left(\Sigma\right)\text{ with probability }1.$ (14)
Proof.
See Appendix A.4
The popular software STAN (Carpenter et al., , 2017) modifies Geyer, ’s (1992) ISE estimator for the parallel chain scenario using an ad-hoc adjustment in the following way. In the one-dimensional setup, for $1\leq m\leq M$, denote
$s_{m}^{2}=\frac{1}{(n-1)}\sum_{t=1}^{n}\left(g(X_{t}^{(m)})-\bar{g}^{(m)}\right)^{2}.$
Further, denote $\widehat{W}=M^{-1}\sum_{m=1}^{M}s_{m}^{2}$ and
$\widehat{B}=\frac{n}{(M-1)}\sum_{m=1}^{M}(\bar{g}^{(m)}-\bar{\bar{g}})^{2}.$
Let $\hat{\gamma}^{(m)}_{i;n}$ denote the lag-$i$ sample autocovariance for the $m^{th}$ chain. Then the lag-$i$ auto-covariance proposed by STAN is
$\hat{\gamma}_{i;n}^{\text{STAN}}=\frac{1}{n}(\widehat{B}-\widehat{W})+\frac{1}{M}\sum_{m=1}^{M}\hat{\gamma}^{(m)}_{i;n}.$
The initial sequence estimators based on these auto-covariance estimators are denoted as $\hat{\sigma}^{2}_{\text{STAN-ISE}}$, and their theoretical properties have not been studied. Similar to before, we can build the covariance-correlation estimator using STAN auto-covariances for the multivariate settings, i.e., the diagonal matrix with STAN-ISE standard deviations can be defined as
$\hat{L}_{\text{STAN-ISE}}=\text{diagonal}\left(\hat{\sigma}_{\text{STAN-ISE}}^{(1)},\hat{\sigma}_{\text{STAN-ISE}}^{(2)},\dots,\hat{\sigma}_{\text{STAN-ISE}}^{(d)}\right)\,.$
The STAN-based multivariate covariance-correlation estimator is then defined as
$\hat{\Sigma}_{\text{STAN-cc}}=\hat{L}_{\text{G-ISE}}\hat{\mathcal{R}}_{\text{BM}}\hat{L}_{\text{G-ISE}}\,.$
The performance of estimators $\Sigma_{\text{G-cc}}$ and $\Sigma_{\text{STAN-cc}}$ are compared for a number of different parallel chains $m=2,4,8$ and $16$ for the vector auto-regressive process of Section 4. We compare the relative Frobenius norm using samples of sizes $10^{3}$ and $10^{5}$ with $100$ replications. Results are in Figure 3. It is clear that although for larger sample sizes $\Sigma_{\text{G-cc}}$ and $\hat{\Sigma}_{\text{STAN-cc}}$ are comparable, for smaller sample sizes, $\hat{\Sigma}_{\text{G-cc}}$ is far more reliable.
Figure 3: Boxplots of estimated relative Frobenious norm for various number of parallel chains for (left) $n=10^{3}$ and (right) $n=10^{5}$.
6 Discussion
Our central goal was to arrive at a version of a multivariate ISE that is computationally efficient enough to be practically usable and that provides reliable finite-sample results. Since the family of multivariate ISE estimators relies on the estimation of autocovariance matrices, this has been traditionally challenging. Our solution provides a practical, efficient, and effective alternative to that of Dai and Jones, ’s (2017) mISE.
SV estimators also require the estimation of sample lag autocovariance matrices and in principle should be computationally involved. The key difference is that the truncation point is determined a priori and not sequentially. In fact, for a fixed truncation point, Heberle and Sattarhoff, (2017) propose an algorithm that also relies on discrete Fourier and fast Fourier transforms to yield an efficient implementation of the SV estimator. However, as is evident from our simulations, our proposed alternative yields superior performance at comparative costs. Recently, Song and Berg, (2023) propose a multivariate version of their MLS estimator, which requires running their MLS estimator for every element of the $\Sigma$ matrix. Given the computational cost evidenced in Section 4 of estimating $d$ diagonals, this estimator is expected to be computationally intensive.
Finally, reversibility of the Markov chain is a critical assumption for employing any initial sequence estimator. It remains an open problem to develop a version of an ISE estimator that is compatible with non-reversible, but stationary Markov chains.
7 Acknowledgements
Dootika Vats is supported by SERB (SPG/2021/001322).
Appendix A Appendix
A.1 Proof of Proposition 1
Theorem 4.
(Dai and Jones, , 2017, Proposition $2$) Let $\Sigma_{m}=-\zeta_{0}+2\sum_{i=0}^{m}Z_{i}$ where $Z_{m}=\zeta_{2m}+\zeta_{2m+1}$ for all $m\geq 0$. Then there exists an integer $m_{0}$ such that
  • 1.
    $\Sigma_{m}$ is positive definite for all $m\geq m_{0}$ and not positive definite for all $m<m_{0}$; and
  • 2.
    the sequence $\{|\Sigma_{m}|:m\geq m_{0}\}$ is positive, increasing, and converges to $|\Sigma|$.
Proof of Proposition 1.
Define $\hat{\Delta}_{i;n}=|\hat{\Sigma}_{i;n}|-|\hat{\Sigma}_{i-1;n}|$ and $\Delta_{i}=|\Sigma_{i}|-|\Sigma_{i-1}|$ for all $i\geq 1$. From ) in Dai and Jones, (2017, Equation (B5)), for all $i$, $\hat{\Delta}_{i;n}\rightarrow\Delta_{i}$ as $n\rightarrow\infty$ with probability $1$. Define the set of events $A_{t}$ for all $t\geq 0$ such that
$A_{t}:=\left\{\lim_{n\rightarrow\infty}\hat{\Delta}_{t;n}=\Delta_{t}\right\}.$
From Dai and Jones, (2017),
$\Pr(A_{t}^{c})=0\,,$ (15)
for all $t\geq 0$. Define event $B=\{s_{n}\leq M\text{ for some }M\in\mathbb{N}\}$. Then by Dai and Jones, (2017, Theorem $2$),
$\Pr(B)=1\Rightarrow\Pr(B^{c})=0$ (16)
By the Boole’s inequality
$\Pr\left((\cup_{t=s_{n}+1}^{\infty}A_{t}^{c})\cup B^{c}\right)\leq\sum_{t=s_{n}+1}^{\infty}\Pr(A_{t}^{c})+\Pr(B^{c})$
$\Rightarrow$ $\Pr\left((\cap_{t=s_{n}+1}^{\infty}A_{t}\cap B)^{c}\right)\leq\sum_{t=s_{n}+1}^{\infty}\Pr(A_{t}^{c})+\Pr(B^{c})$
[…]
$\Rightarrow$ $\Pr(\cap_{t=s_{n}+1}^{\infty}A_{t}\cap B)=1$
$\Rightarrow$ $\Pr\left(\lim_{n\rightarrow\infty}\hat{\Delta}_{i;n}=\Delta_{i}\quad\text{ for all }i\geq s_{n}+1\right)=1.$ (17)
Now
$\Pr(t_{n}\rightarrow\infty\text{ as }n\rightarrow\infty)$
$=$ $\Pr\left(|\hat{\Sigma}_{i;n}|>|\hat{\Sigma}_{i-1;n}|\text{ as }n\rightarrow\infty\text{ for all }i\geq s_{n}+1\right)$
$=$ $\Pr\left(|\hat{\Sigma}_{i;n}|-|\hat{\Sigma}_{i-1;n}|>0\text{ as }n\rightarrow\infty\text{ for all }i\geq s_{n}+1\right)$
$\geq$ $\Pr\left(\hat{\Delta}_{i;n}\rightarrow\Delta_{i}\text{ as }n\rightarrow\infty\text{ for all }i\geq s_{n}+1\right)$
$=$ $1$
A.2 Proof of Theorem 2
Proof.
We show each part separately.
  • 1.
    Note that
    $|\hat{\Sigma}_{\text{cc}}|$ $=|\hat{L}_{\text{ISE}}||\hat{L}^{\top}_{\text{ISE}}||\hat{\mathcal{R}}_{\text{BM}}|$
    $=\left(\prod_{i=1}^{d}\left(\hat{\sigma}_{\text{ISE}}^{(i)}\right)^{2}\right)|\hat{\mathcal{R}}_{\text{BM}}|.$
    Under the Assumption 1 and 2 and from Vats et al., (2019) Theorem $2$, $\hat{\Sigma}_{\text{BM}}\rightarrow\Sigma$ as $n\rightarrow\infty$ with probability $1$. Also define $\hat{a}_{ij;\text{BM}}$ to be the $(i,j)^{th}$ elements of $\hat{\Sigma}_{\text{BM}}$. Consequently, $\hat{a}_{ij;\text{BM}}\rightarrow a_{ij}$ as $n\rightarrow\infty$ with probability $1$. For $\hat{a}_{ii;\text{BM}}>0$ for all $i$ and by continuous mapping theorem, with probability 1
    $\frac{\hat{a}_{ij;\text{BM}}}{\sqrt{\hat{a}_{ii;\text{BM}}\ \hat{a}_{jj;\text{BM}}}}$ $\rightarrow\frac{a_{ij}}{\sqrt{a_{ii}\ a_{jj}}}\ \text{as }n\rightarrow\infty$
    $\Rightarrow\hat{r}_{ij}$ $\rightarrow r_{ij}\ \text{as }n\rightarrow\infty$
    where $\hat{r}_{ij}$ and $r_{ij}$ are $(i,j)^{th}$ elements of ^$\hat{\mathcal{R}}_{\text{BM}}$ and $\mathcal{R}$ respectively. Hence as $n\rightarrow\infty$
    ^$\hat{\mathcal{R}}_{\text{BM}}\rightarrow\mathcal{R}\text{ with probability }1.$ (18)
    By Vats et al., (2018, Theorem $3$), the eigenvalues of ^$\hat{\mathcal{R}}_{\text{BM}}$ are strongly consistent to eigenvalues of $\mathcal{R}$. This implies as $n\rightarrow\infty$ with probability $1$
    |^$|\hat{\mathcal{R}}_{\text{BM}}|\rightarrow|\mathcal{R}|.$ (19)
    Additionally, from Geyer, (1992, Theorem $3$) as $n\rightarrow\infty$ with probability $1$ for $i=1,2,\cdots,d$
    $\liminf_{n\rightarrow\infty}\left(\hat{\sigma}_{\text{ISE}}^{(i)}\right)^{2}\geq\left(\sigma^{(i)}\right)^{2}.$ (20)
    Hence, from (19) and (20) with probability $1$
    $\liminf_{n\rightarrow\infty}|\hat{\Sigma}_{\text{cc}}|$ $=\left(\prod_{i=1}^{d}\liminf_{n\rightarrow\infty}\left(\hat{\sigma}_{\text{ISE}}^{(i)}\right)^{2}\right)\lim_{n\rightarrow\infty}|\hat{\mathcal{R}}_{\text{BM}}|$
    $\geq\left(\prod_{i=1}^{d}\left(\sigma^{(i)}\right)^{2}\right)|\mathcal{R}|$
    $=|L\mathcal{R}L|$
    $=|\Sigma|.$
  • 2.
    From (18), as $n\rightarrow\infty$
[…]
    Hence with probability $1$
    $\liminf_{n\rightarrow\infty}|\hat{a}_{ij}|$ $=\liminf_{n\rightarrow\infty}\left(\hat{\sigma}_{\text{ISE}}^{(i)}|\hat{r}_{ij}|\hat{\sigma}_{\text{ISE}}^{(j)}\right)$
    $\geq\sigma^{(i)}|r_{ij}|\sigma^{(j)}$
    $=|a_{ij}|.$
A.3 Proof of Theorem 3
Proof.
First, we show that globally centered auto-covariance matrices, averaged over all the parallel chains are strongly consistent for the true auto-covariance matrix. Hence for all $t\geq 0$ as $n\rightarrow\infty$
$\hat{\gamma}_{t;n}^{\text{G}}$ $=\frac{1}{M}\sum_{s=1}^{M}\hat{\gamma}^{(s)}_{t;n}$
$=\frac{1}{M}\sum_{s=1}^{M}\frac{1}{n-t}\sum_{i=1}^{n-t}\left(g(X_{t}^{(s)})-\bar{\bar{g}}\right)\left(g(X_{t+i}^{(s)})-\bar{\bar{g}}\right)$
$=\frac{1}{M}\sum_{s=1}^{M}\frac{1}{n-t}\sum_{i=1}^{n-t}\left[g(X_{t}^{(s)})g(X_{t+i}^{(s)})-g(X_{t}^{(s)})\bar{\bar{g}}-\bar{\bar{g}}g(X_{t+i}^{(s)})+\bar{\bar{g}}^{2}\right]$
[…]
Now the globally-centered univariate ISE is defined as
$\hat{\sigma}_{\text{G-ISE}}^{2}=-\hat{\gamma}_{0;n}^{\text{G}}+2\sum_{i=0}^{k_{n}}(\hat{\gamma}_{2t;n}^{\text{G}}+\hat{\gamma}_{2t+1;n}^{\text{G}})$
where $k_{n}$ is the stopping time as defined in (11). Note that for every $\epsilon>0$, there exists a stopping time $m_{\epsilon}>0$ such that
$2\sum_{i=m_{\epsilon}+1}^{\infty}(\gamma_{2t}+\gamma_{2t+1})$ $<\epsilon/2$
$\Rightarrow-\gamma_{0}+2\sum_{i=0}^{m_{\epsilon}}(\gamma_{2t}+\gamma_{2t+1})$ $>\sigma^{2}-\epsilon/2.$ (21)
Again, for every $\epsilon>0$, there exists a sample size $n_{\epsilon}>0$ such that for all $n>n_{\epsilon}$, and $t\in\{0,1,\ldots,m_{\epsilon}\}$, $(\hat{\gamma}_{2t;n}^{\text{G}}+\hat{\gamma}_{2t+1;n}^{\text{G}})>0$ and
$-\gamma_{0}+2\sum_{i=0}^{m_{\epsilon}}(\gamma_{2t}+\gamma_{2t+1})-\left(-\hat{\gamma}_{0;n}^{\text{G}}+2\sum_{i=0}^{m_{\epsilon}}(\hat{\gamma}_{2t;n}^{\text{G}}+\hat{\gamma}_{2t+1;n}^{\text{G}})\right)<\epsilon/2$
$\Rightarrow$ $-\hat{\gamma}_{0;n}^{\text{G}}+2\sum_{i=0}^{m_{\epsilon}}(\hat{\gamma}_{2t;n}^{\text{G}}+\hat{\gamma}_{2t+1;n}^{\text{G}})>-\gamma_{0}+2\sum_{i=0}^{m_{\epsilon}}(\gamma_{2t}+\gamma_{2t+1})-\epsilon/2$
[…]
$\Rightarrow$ $\inf_{n>n_{\epsilon}}\left(-\hat{\gamma}_{0;n}^{\text{G}}+2\sum_{i=0}^{m_{\epsilon}}(\hat{\gamma}_{2t;n}^{\text{G}}+\hat{\gamma}_{2t+1;n}^{\text{G}})\right)>\sigma^{2}-\epsilon.$
Note that as $\epsilon\rightarrow 0$, $n_{\epsilon}\rightarrow\infty$ and the stopping time $m_{\epsilon}\rightarrow\infty$. Also by definition, for all $n>n_{\epsilon}$, and $t\in\{0,1,\ldots,m_{\epsilon}\}$, $(\hat{\gamma}_{2t;n}^{\text{G}}+\hat{\gamma}_{2t+1;n}^{\text{G}})>0$ and it implies that $k_{n}\geq m_{\epsilon}$. Hence
$\liminf_{n\rightarrow\infty}\left(-\hat{\gamma}_{0;n}^{\text{G}}+2\sum_{i=0}^{m_{\epsilon}}(\hat{\gamma}_{2t;n}^{\text{G}}+\hat{\gamma}_{2t+1;n}^{\text{G}})\right)\geq\sigma^{2}$
$\liminf_{n\rightarrow\infty}\left(-\hat{\gamma}_{0;n}^{\text{G}}+2\sum_{i=0}^{k_{n}}(\hat{\gamma}_{2t;n}^{\text{G}}+\hat{\gamma}_{2t+1;n}^{\text{G}})\right)\geq\sigma^{2}\text{ (since }k_{n}\geq m_{\epsilon})$
$\Rightarrow$ $\liminf_{n\rightarrow\infty}\hat{\sigma}^{2}_{\text{G-ISE}}\geq\sigma^{2}.$ (22)
A.4 Proof of Corollary 1
Proof.
Let $\{X_{t}^{(s)}\}_{t\geq 1}$ be a set of parallel Markov chains for $s=1,2,\ldots,M$. By Theorem 3 as $n\rightarrow\infty$ and for all $j=1,2,\ldots,M$
$\liminf_{n\rightarrow\infty}\left(\hat{\sigma}_{\text{G-ISE}}^{(j)}\right)^{2}\geq\left(\sigma^{(j)}\right)^{2}.$
Hence
$\liminf_{n\rightarrow\infty}\text{det}\left(\hat{L}_{\text{G-ISE}}\right)=\prod_{j=1}^{d}\hat{\sigma}_{\text{G-ISE}}^{(j)}\geq\text{det}\left(L\right).$ (23)
Also, under Assumptions 1 and 2, by Gupta and Vats, (2020, Theorem $2$),
$\lim_{n\rightarrow\infty}\hat{\Sigma}_{\text{G-BM}}$ $=\Sigma\text{ with probability }1$
$\Rightarrow\lim_{n\rightarrow\infty}\hat{\mathcal{R}}_{\text{G-BM}}$ = with probability 1.absent $=\mathcal{R}\text{ with probability }1.$
Hence, by Vats et al., (2018, Theorem $3$), the eigenvalues of ^$\hat{\mathcal{R}}_{\text{G-BM}}$ are strongly consistent to the eigenvalues of $\mathcal{R}$. This implies as $n\rightarrow\infty$
$\lim_{n\rightarrow\infty}\text{det}\left(\hat{\mathcal{R}}_{\text{G-BM}}\right)=\text{det}\left(\mathcal{R}\right)\text{ with probability }1.$ (24)
By (23) and (24)
$\liminf_{n\rightarrow\infty}\text{det}\left(\hat{\Sigma}_{\text{G-cc}}\right)=\liminf_{n\rightarrow\infty}\text{det}\left(\hat{L}_{\text{G-ISE}}\hat{\mathcal{R}}_{\text{G-BM}}\hat{L}_{\text{G-ISE}}\right)\geq\text{det}\left(L\mathcal{R}L\right)=\text{det}\left(\Sigma\right).$